# Learning Constrained Markov Decision Processes With Non-stationary Rewards and Constraints

### v1

@@ 2.1 Constrained Markov decision processes
$C_{G}\coloneqq{\min_{g\in[0,1]^{|X\times A|}}}\underset{t\in[T]}{\sum}\underset{i\in[m]}{\max}\lVert\mathbb{E}[g_{t,i}]-g\rVert_{1}.$ (2)
We let $g^{\circ}\in[0,1]^{|X\times A|}$ be the constraint cost vector that attains the minimum in the definition of $C_{G}$. Finally, we introduce the total adversarial corruption $C$, which is defined as $C\coloneqq\max\{C_{G},C_{r}\}$. [^3] In the rest of the paper, for ease of presentation, we derive all our results in terms of the total adversarial corruption $C$. We remark that such results can be easily extended to account for $C_{r}$ and $C_{G}$ separately.
@@ 3.1 NS-SOPS: non-stationary safe optimistic policy search
Intuitively, the first term in the expressions above is derived from Azuma-Hoeffding inequality, the second term allows to deal with the non-stationarity of rewards and constraint costs, while the third term is needed to bound how much the average reward vector $\overline{r}$ and the average constraint costs $[\overline{G}]_{i}$ differ from their “fictitious” non-corrupted counterparts $r^{\circ}$ and $g^{\circ}$, respectively.
@@ 3.2 Theoretical guarantees of NS-SOPS
$V_{T}=\mathcal{O}\left(L|X||A|\sqrt{T\ln\left(\nicefrac{{mT|X||A|}}{{\delta}}\right)}+\ln(T)|X||A|C\right).$
@@
$R_{T}=\mathcal{O}\left(L|X||A|\sqrt{T\ln\left(\nicefrac{{T|X||A|}}{{\delta}}\right)}+\ln(T)|X||A|C\right).$
@@ 4.1 Lag-FTRL: Lagrangified FTRL
At a high level, the Lagrangified follow-the-regularized-leader (Lag-FTRL for short) algorithm works by instantiating several different instances of Algorithm 2, with each instance $\texttt{Alg}^{j}$ being run for a different “guess” of the (unknown) adversarial corruption value $C$. The algorithm plays the role of a master by choosing which instance $\texttt{Alg}^{j}$ to use at each episode. The selection is done by employing an FTRL approach with a suitable log-barrier regularization. In particular, at each episode $t\in[T]$, by letting $\texttt{Alg}^{j_{t}}$ be the selected instance, the Lag-FTRL algorithm employs the policy $\pi_{t}^{j_{i}}$ prescribed by $\texttt{Alg}^{j_{t}}$ and provides the observed feedback to instance $\texttt{Alg}^{j_{t}}$ only.
@@
3:$\gamma\leftarrow\sqrt{\nicefrac{{\ln(M/\delta)}}{{TM}}}$, $\eta\leftarrow\frac{1}{2Lm\Lambda}\sqrt{\nicefrac{{\ln(T)}}{{T}}}$
@@
5:     $\texttt{Alg}^{j}\leftarrow$ Algorithm 2 with $C=2^{j}$
@@
10:     Choose policy $\pi_{t}^{j_{t}}$ in Algorithm 1 and observe feedback from interaction
@@
14:     $w_{t+1}\leftarrow\operatorname*{arg\,min}_{w\in\Delta_{M}}\,w^{\top}\sum_{\tau\in[t]}\ell_{t}+\frac{1}{\eta}\sum_{j\in[M]}\frac{1}{w_{j}}$
The pseudocode of the Lag-FTRL algorithm is provided in Algorithm 3. At Line 5, it instantiates $M\coloneqq\lceil\log_{2}T\rceil$ instances of Algorithm 2, with each instance $\texttt{Alg}^{j}$, for $j\in[M]$, receiving as input a “guess” on the adversarial corruption $C=2^{j}$. The algorithm assigns weights defining a probability distribution to instances $\texttt{Alg}^{j}$, with $w_{t,j}\in[0,1]$ denoting the weight of instance $\texttt{Alg}^{j}$ at episode $t\in[T]$. We denote by $w_{t}\in\Delta_{M}$ the weight vector at episode $t$, with $\Delta_{M}$ being the $M$-dimensional simplex. At the first episode, all the weights $w_{1,j}$ are initialized to the value $\nicefrac{{1}}{{M}}$ (Line 6). Then, at each episode $t\in[T]$, the algorithm samples an instance index $j_{t}\in[M]$ according to the probability distribution defined by the weight vector $w_{t}$ (Line 8), and it employs the policy $\pi_{t}^{j_{t}}$ prescribed by $\texttt{Alg}^{j_{t}}$ (Line 9). The algorithm observes the feedback from the interaction described in Algorithm 1 and it sends such a feedback to instance $\texttt{Alg}^{j_{t}}$ (Line 11). Then, at Line 13, the algorithm builds an optimistic loss estimator to be fed into each instance $\texttt{Alg}^{j}$. In particular, at episode $t\in[T]$ and for every $j\in[M]$, the optimistic loss estimator is defined as:
@@
where $\gamma$ is a suitably-defined implicit exploration factor, $(x_{k}^{t},a_{k}^{t})$ is the state-action pair visited at layer $k$ during episode $t$, $\Lambda$ is a suitably-defined upper bound on the optimal values of Lagrangian multipliers, [^5] Notice that, in the definition of $\Lambda$, $\rho$ is the feasibility parameter of Program (6) for the reward vector $\overline{r}$, the constraint cost matrix $\overline{G}$, and the threshold vector $\alpha$. In order to compute $\Lambda$, Algorithm 3 needs knowledge of $\rho$. Nevertheless, our results continue to hold even if Algorithm 3 is only given access to a lower bound on $\rho$. $\widehat{G}_{t}^{j}$ is the matrix of empirical constraint costs built by the instance $\texttt{Alg}^{j}$ of Algorithm 2 at episode $t$, while $\widehat{q}_{t}^{j}$ is the occupancy measure computed by instance $\texttt{Alg}^{j}$ of Algorithm 2 at $t$. Finally, the algorithm updates the weight vector according to an FTRL update with a suitable log-barrier regularization (see Line 14 for the definition of the update).
@@ 4.2 Theoretical guarantees of Lag-FTRL
$V_{T}=\mathcal{O}\left(m^{2}L^{2}|X||A|\sqrt{T\ln\left(\nicefrac{{mT|X||A|}}{{\delta}}\right)}\ln(T)+m^{2}L\ln(T)|X||A|C\right).$
@@
If Program (6) instantiated with $\overline{r}$, $\overline{G}$ and $\alpha$ satisfies Slater’s condition (Condition 1), then, given any $\delta\in(0,1)$, Algorithm 3 attains the following bound with probability at least $1-26\delta$:
$R_{T}=\mathcal{O}\left(m^{2}L^{2}|X||A|\sqrt{T\ln\left(\nicefrac{{mT|X||A|}}{{\delta}}\right)}\ln(T)+m^{2}L\ln(T)|X||A|C\right).$
@@ Appendix A Related works
The literature on online learning problems (Cesa-Bianchi and Lugosi, 2006) in MDPs is wide (see (Auer et al., 2008; Even-Dar et al., 2009; Neu et al., 2010) for some results on the topic). In such settings, two types of feedback are usually studied: in the full-information feedback model, the entire loss function is observed after the learner’s choice, while in the bandit feedback model, the learner only observes the loss due to the chosen action. Azar et al. (2017) study the problem of optimal exploration in episodic MDPs with unknown transitions and stochastic losses when the feedback is bandit. The authors present an algorithm whose regret upper bound is $\tilde{\mathcal{O}}(\sqrt{T})$, thus matching the lower bound for this class of MDPs and improving the previous result by Auer et al. (2008). Finally, online learning in MDPs has been studied form the configurator perspective, namely, the agent whose actions are different possible transition functions of the underlying environment (see, (Maran et al., 2024)).
@@
The literature on non-stationary MDPs encompasses both works on non-stationary rewards and non-stationary transitions. As concerns the first research line, Rosenberg and Mansour (2019b) study the online learning problem in episodic MDPs with adversarial losses and unknown transitions when the feedback is full information. The authors present an online algorithm exploiting entropic regularization and providing a regret upper bound of $\tilde{\mathcal{O}}(\sqrt{T})$. The same setting is investigated by Rosenberg and Mansour (2019a) when the feedback is bandit. In such a case, the authors provide a regret upper bound of the order of $\tilde{\mathcal{O}}(T^{3/4})$, which is improved by Jin et al. (2020) by providing an algorithm that achieves in the same setting a regret upper bound of $\tilde{\mathcal{O}}(\sqrt{T})$. Furthermore, Bacchiocchi et al. (2023) study adversarial MDPs where the bandit feedback depends on a different agent. Related to the non-stationarity of the transitions , Wei et al. (2022) study MDPs with adversarial corruption on transition functions and rewards, reaching a regret upper bound of order $\widetilde{\mathcal{O}}(\sqrt{T}+C)$ (where $C$ is the amount of adversarial corruption) with respect to the optimal policy of the non-corrupted MDP. Finally, Jin et al. (2024) is the first to study completely adversarial MDPs with changing transition functions, providing a $\tilde{\mathcal{O}}(\sqrt{T}+C)$ regret bounds, where $C$ is a corruption measure of the adversarially changing transition functions.
@@
Online Learning In MDPs with constraints is generally studied when the constraints are selected stochastically (see (Efroni et al., 2020)). Precisely, Zheng and Ratliff (2020) deal with episodic CMDPs with stochastic losses and constraints, where the transition probabilities are known and the feedback is bandit. The regret upper bound of their algorithm is of the order of $\tilde{\mathcal{O}}(T^{3/4})$, while the cumulative constraint violation is guaranteed to be below a threshold with a given probability. Wei et al. (2018) deal with adversarial losses and stochastic constraints, assuming the transition probabilities are known and the feedback is full information. The authors present an algorithm that guarantees an upper bound of the order of $\tilde{\mathcal{O}}(\sqrt{T})$ on both regret and constraint violation. Bai et al. (2020) provide the first algorithm that achieves sublinear regret when the transition probabilities are unknown, assuming that the rewards are deterministic and the constraints are stochastic with a particular structure. Efroni et al. (2020) propose two approaches to deal with the exploration-exploitation dilemma in episodic CMDPs. These approaches guarantee sublinear regret and constraint violation when transition probabilities, rewards, and constraints are unknown and stochastic, while the feedback is bandit. Qiu et al. (2020) provide a primal-dual approach based on optimism in the face of uncertainty. This work shows the effectiveness of such an approach when dealing with episodic CMDPs with adversarial losses and stochastic constraints, achieving both sublinear regret and constraint violation with full-information feedback. Stradi et al. (2024) is the first work to tackle CMDPs with adversarial losses and bandit feedback. They propose an algorithm which achieves sublinear regret and sublinear positive constraints violations, assuming that the constraints are stochastic. Germano et al. (2023) are the first to study CMDPs with adversarial constraints. Given the well-known impossibility result to learn with adversarial constraints, they propose an algorithm that attains sublinear violation (with cancellations allowed) and a fraction of the optimal reward when the feedback is full. Finally, Ding and Lavaei (2023) and Wei et al. (2023) consider the case in which rewards and constraints are non-stationary, assuming that their variation is bounded, as in our work. Nevertheless, our settings differ in multiple aspects. First of all, we consider positive constraints violations, while the aforementioned works allow the cancellations in their definition. We consider a static regret adversarial baseline, while Ding and Lavaei (2023) and Wei et al. (2023) consider the stronger baseline of dynamic regret. Nevertheless, our bounds are not comparable, since we achieve linear regret and violations only in the worst case scenario in which $C=T$, while a sublinear corruption would lead to linear dynamic regret in their work. Finally, we do not make any assumption on the number of episodes, while both the regret and violations bounds presented in Wei et al. (2023) hold only for large $T$. Finally, Bacchiocchi et al. (2024) study a generalization of stochastic CMDPs with partial observability on the constraints.

### v2

@@ 2.1 Constrained Markov decision processes
$C_{G}\coloneqq{\min_{G\in[0,1]^{|X\times A|\times m}}}\underset{t\in[T]}{\sum}\underset{i\in[m]}{\max}\lVert\mathbb{E}[g_{t,i}]-g_{i}\rVert_{1},$ (2)
where $g_{i}$ is the $i$-th component of $G$. We let $G^{\circ}\in[0,1]^{|X\times A|\times m}$ be the constraint cost matrix that attains the minimum in the definition of $C_{G}$. Finally, we introduce the total adversarial corruption $C$, which is defined as $C\coloneqq\max\{C_{G},C_{r}\}$.
@@ 3.1 NS-SOPS: non-stationary safe optimistic policy search
Intuitively, the first term in the expressions above is derived from Azuma-Hoeffding inequality, the second term allows to deal with the non-stationarity of rewards and constraint costs, while the third term is needed to bound how much the average reward vector $\overline{r}$ and the average constraint costs $[\overline{G}]_{i}$ differ from their “fictitious” non-corrupted counterparts $r^{\circ}$ and $[G^{\circ}]_{i}$, respectively.
@@ 3.2 Theoretical guarantees of NS-SOPS
$V_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\nicefrac{{mT|X||A|}}{{\delta}}\right)}+\ln(T)|X||A|C\right).$
@@
$R_{T}=\mathcal{O}\left(L|X|\sqrt{|A|T\ln\left(\nicefrac{{T|X||A|}}{{\delta}}\right)}+\ln(T)|X||A|C\right).$
@@ 4.1 Lag-FTRL: Lagrangified FTRL
At a high level, the Lagrangified follow-the-regularized-leader (Lag-FTRL for short) algorithm works by instantiating several different instances of Algorithm 2, suitably stabilized (see section H), with each instance $\texttt{Alg}^{j}$ being run for a different “guess” of the (unknown) adversarial corruption value $C$. The algorithm plays the role of a master by choosing which instance $\texttt{Alg}^{j}$ to use at each episode. The selection is done by employing an FTRL approach with a suitable log-barrier regularization. In particular, at each episode $t\in[T]$, by letting $\texttt{Alg}^{j_{t}}$ be the selected instance, the Lag-FTRL algorithm employs the policy $\pi_{t}^{j_{i}}$ prescribed by $\texttt{Alg}^{j_{t}}$ and provides the observed feedback to instance $\texttt{Alg}^{j_{t}}$ only.
@@
3:$\gamma\leftarrow\sqrt{\nicefrac{{\ln(M/\delta)}}{{TM}}}$,  $\eta\leftarrow$$\frac{1}{2\Lambda m\left(\sqrt{\beta_{1}T}+\beta_{2}+\beta_{5}+\sqrt{\beta_{4}T}\right)}$
@@
5:     $\texttt{Alg}^{j}\leftarrow$ stabilized Algorithm 2 with $C=2^{j}$
@@
10:     Choose policy $\pi_{t}^{j_{t}}$ in Algorithm 1 and observe ……feedback from interaction
@@
14:         Build $b_{t,j}$ as in Equation (9)
15:     $w_{t+1}\leftarrow\operatorname*{arg\,min}_{\underset{w_{j}\geq\nicefrac{{1}}{{T}}}{w\in\Delta_{M},}}\,w^{\top}\sum_{\tau\in[t]}(\ell_{t}-b_{t})+\frac{1}{\eta}\sum_{j\in[M]}\ln\frac{1}{w_{j}}$
The pseudocode of the Lag-FTRL algorithm is provided in Algorithm 3. At Line 5, it instantiates $M\coloneqq\lceil\log_{2}T\rceil$ instances of Algorithm 2, with each instance $\texttt{Alg}^{j}$, for $j\in[M]$, receiving as input a “guess” on the adversarial corruption $C=2^{j}$. Notice that, to every instance of Algorithm 2, a standard doubling trick and a stabilization procedure is applied (see Algorithm 4 for additional details). This modification to Algorithm 2 is necessary to guarantee that each instance $j$ attains a regret and positive cumulative constraints violation which smoothly degrade with $\nu_{T,j}=\nicefrac{{1}}{{\min_{t\in[T]}w_{t,j}}}$ and linearly in $C$, when employed by the master algorithm. The algorithm assigns weights defining a probability distribution to instances $\texttt{Alg}^{j}$, with $w_{t,j}\in[0,1]$ denoting the weight of instance $\texttt{Alg}^{j}$ at episode $t\in[T]$. We denote by $w_{t}\in\Delta_{M}$ the weight vector at episode $t$, with $\Delta_{M}$ being the $M$-dimensional simplex. At the first episode, all the weights $w_{1,j}$ are initialized to the value $\nicefrac{{1}}{{M}}$ (Line 6). Then, at each episode $t\in[T]$, the algorithm samples an instance index $j_{t}\in[M]$ according to the probability distribution defined by the weight vector $w_{t}$ (Line 8), and it employs the policy $\pi_{t}^{j_{t}}$ prescribed by $\texttt{Alg}^{j_{t}}$ (Line 9). The algorithm observes the feedback from the interaction described in Algorithm 1 and it sends such a feedback to instance $\texttt{Alg}^{j_{t}}$ (Line 11). Then, at Line 13, the algorithm builds an optimistic loss estimator to be fed into each instance $\texttt{Alg}^{j}$. In particular, at episode $t\in[T]$ and for every $j\in[M]$, the optimistic loss estimator is defined as:
@@
where $\gamma$ is a suitably-defined implicit exploration factor, $(x_{k}^{t},a_{k}^{t})$ is the state-action pair visited at layer $k$ during episode $t$, $\Lambda$ is a suitably-defined upper bound on the optimal values of Lagrangian multipliers, [^4] Notice that, in the definition of $\Lambda$, $\rho$ is the feasibility parameter of Program (6) for the reward vector $\overline{r}$, the constraint cost matrix $\overline{G}$, and the threshold vector $\alpha$. In order to compute $\Lambda$, Algorithm 3 needs knowledge of $\rho$. Nevertheless, our results continue to hold even if Algorithm 3 is only given access to a lower bound on $\rho$. $\widehat{G}_{t}^{j}$ is the matrix of empirical constraint costs built by the instance $\texttt{Alg}^{j}$ of Algorithm 2 at episode $t$, while $\widehat{q}_{t}^{j}$ is the occupancy measure computed by instance $\texttt{Alg}^{j}$ of Algorithm 2 at $t$. Finally, the algorithm updates the weight vector according to an FTRL update on a cut decision space with a suitable log-barrier regularization and a bonus term $b_{t}$ defined as:
$b_{t,j}\coloneq\left(\left(m\Lambda\beta_{5}+\beta_{2}\right)+\left(\sqrt{\beta_{1}}+m\Lambda\sqrt{\beta_{4}}\right)\sqrt{T}\right)(\nu_{t,j}-\nu_{t-1,j}),,$ (9)
where $\nu_{t,j}=\max_{\tau\leq t}\frac{1}{w_{\tau,j}}$ and the parameters $\beta$ are linked to the performance of Algorithm 2 (see Line 14 and Section F.2.1 for additional details). See Line 15 for the complete definition of the update. The bonus term purpose is to balance out the term related to the difference between the performance of Algorithm 2 updated at each episode and the performance of its stabilized version, which works under the condition imposed by the master algorithm.
@@ 4.2 Theoretical guarantees of Lag-FTRL
If Program (6) instantiated with $\overline{r}$, $\overline{G}$ and $\alpha$ satisfies Slater’s condition (Condition 1), then, given any $\delta\in(0,1)$, Algorithm 3 attains the following bound with probability at least $1-34\delta$:
$|\displaystyle V_{T}=\mathcal{O}\bigg{(}m^{2}L^{2}|X|$$\sqrt{|A|T\log\left(\nicefrac{{mT|X||A|}}{{\delta}}\right)}\log(T)^{2}$
$+m^{2}L|X|^{2}|A|^{2}\log(T)^{3}\log\left(\nicefrac{{\log(T)}}{{\delta}}\right)+m^{2}L\log(T)^{2}|X||A|C\bigg{)}.$
@@
$|\displaystyle R_{T}=\mathcal{O}\bigg{(}m^{2}L^{2}|X|$$\sqrt{|A|T\log\left(\nicefrac{{mT|X||A|}}{{\delta}}\right)}\log(T)^{2}$
$+m^{2}L|X|^{2}|A|^{2}\log(T)^{3}\log\left(\nicefrac{{\log(T)}}{{\delta}}\right)+m^{2}L\log(T)^{2}|X||A|C\bigg{)}.$
@@
  • In Appendix H we provide the results related to stability a corruption-robustness.
@@ Appendix A Related works
The literature on online learning problems (Cesa-Bianchi and Lugosi, 2006) in MDPs is wide (see (Auer et al., 2008; Even-Dar et al., 2009; Neu et al., 2010) for some initial results on the topic). In such settings, two types of feedback are usually studied: in the full-information feedback model, the entire loss function is observed after the learner’s choice, while in the bandit feedback model, the learner only observes the loss due to the chosen action. Azar et al. (2017) study the problem of optimal exploration in episodic MDPs with unknown transitions and stochastic losses when the feedback is bandit. The authors present an algorithm whose regret upper bound is $\tilde{\mathcal{O}}(\sqrt{T})$, thus matching the lower bound for this class of MDPs and improving the previous result by Auer et al. (2008).
@@
The literature on non-stationary MDPs encompasses both works on non-stationary rewards and non-stationary transitions. As concerns the first research line, Rosenberg and Mansour (2019b) study the online learning problem in episodic MDPs with adversarial losses and unknown transitions when the feedback is full information. The authors present an online algorithm exploiting entropic regularization and providing a regret upper bound of $\tilde{\mathcal{O}}(\sqrt{T})$. The same setting is investigated by Rosenberg and Mansour (2019a) when the feedback is bandit. In such a case, the authors provide a regret upper bound of the order of $\tilde{\mathcal{O}}(T^{3/4})$, which is improved by Jin et al. (2020) by providing an algorithm that achieves in the same setting a regret upper bound of $\tilde{\mathcal{O}}(\sqrt{T})$. Related to the non-stationarity of the transitions , Wei et al. (2022) study MDPs with adversarial corruption on transition functions and rewards, reaching a regret upper bound of order $\widetilde{\mathcal{O}}(\sqrt{T}+C)$ (where $C$ is the amount of adversarial corruption) with respect to the optimal policy of the non-corrupted MDP . Finally, Jin et al. (2024) is the first to study completely adversarial MDPs with changing transition functions, providing a $\tilde{\mathcal{O}}(\sqrt{T}+C)$ regret bounds, where $C$ is a corruption measure of the adversarially changing transition functions.
@@
Online Learning In MDPs with constraints is generally studied when the constraints are selected stochastically. Precisely, Zheng and Ratliff (2020) deal with episodic CMDPs with stochastic losses and constraints, where the transition probabilities are known and the feedback is bandit. The regret upper bound of their algorithm is of the order of $\tilde{\mathcal{O}}(T^{3/4})$, while the cumulative constraint violation is guaranteed to be below a threshold with a given probability. Wei et al. (2018) deal with adversarial losses and stochastic constraints, assuming the transition probabilities are known and the feedback is full information. The authors present an algorithm that guarantees an upper bound of the order of $\tilde{\mathcal{O}}(\sqrt{T})$ on both regret and constraint violation. Bai et al. (2020) provide the first algorithm that achieves sublinear regret when the transition probabilities are unknown, assuming that the rewards are deterministic and the constraints are stochastic with a particular structure. Efroni et al. (2020) propose two approaches to deal with the exploration-exploitation dilemma in episodic CMDPs. These approaches guarantee sublinear regret and constraint violation when transition probabilities, rewards, and constraints are unknown and stochastic, while the feedback is bandit. Qiu et al. (2020) provide a primal-dual approach based on optimism in the face of uncertainty. This work shows the effectiveness of such an approach when dealing with episodic CMDPs with adversarial losses and stochastic constraints, achieving both sublinear regret and constraint violation with full-information feedback. Stradi et al. (2024) is the first work to tackle CMDPs with adversarial losses and bandit feedback. They propose an algorithm which achieves sublinear regret and sublinear positive constraints violations, assuming that the constraints are stochastic. Germano et al. (2023) are the first to study CMDPs with adversarial constraints. Given the well-known impossibility result to learn with adversarial constraints, they propose an algorithm that attains sublinear violation (with cancellations allowed) and a fraction of the optimal reward when the feedback is full. Finally, Ding and Lavaei (2023) and Wei et al. (2023) consider the case in which rewards and constraints are non-stationary, assuming that their variation is bounded, as in our work. Nevertheless, our settings differ in multiple aspects. First of all, we consider positive constraints violations, while the aforementioned works allow the cancellations in their definition. We consider a static regret adversarial baseline, while Ding and Lavaei (2023) and Wei et al. (2023) consider the stronger baseline of dynamic regret. Nevertheless, our bounds are not comparable, since we achieve linear regret and violations only in the worst case scenario in which $C=T$, while a sublinear corruption would lead to linear dynamic regret in their work. Finally, we do not make any assumption on the number of episodes, while both the regret and violations bounds presented in Wei et al. (2023) hold only for large $T$.